OpenStudy (anonymous):

SOLVE FOR x -9x+9=-7(x-1)

OpenStudy (anonymous):

@ktnguyen1

OpenStudy (anonymous):

-9x+9=-7x+7 -9x+7x=7-9 -2x=-2 x=1

OpenStudy (anonymous):

9(X-4)= 2x-1 @ktnguyen1

OpenStudy (anonymous):

9x-36=2x-1 9x-2x=36-1 7x=35 x=35/7 x=5

OpenStudy (anonymous):